# Hypergraph Characterization of Fusion Rings

### v1

@@ 5. Low Rank Fusion Rings
$7$ $(1,7)$, $(2,7)$, $(3,7)$ $(1,2,3)$, $(1,2,4)$, $(1,3,5)$, $(1,6,7)$, $(2,3,6)$, $(2,5,7)$, $(3,4,7)$, $(4,5,6)$ $\operatorname{Fib}\boxtimes\operatorname{Sem}^{2}$[21]
$7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(6,3)$, $(6,7)$, $(7,3)$ $(1,2,3)$, $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(3,4,5)$, $(4,6,7)$, $(5,6,7)$
$4$ $(2,4)$, $(3,1)$, $(3,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,4)$, $(2,3,7)$, $(2,5,6)$, $(3,4,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSO}(5)_{6}$[21, 6]
$5$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,6)$, $(2,7)$, $(4,3)$, $(4,6)$, $(5,3)$, $(6,3)$, $(6,7)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
$6$, $7$ $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(6,7)$, $(7,6)$ $(1,2,3)$, $(1,2,6)$, $(1,2,7)$, $(1,4,6)$, $(1,5,7)$, $(2,4,7)$, $(2,5,6)$, $(3,4,5)$, $(3,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{6}$[21, 6]
$6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
$6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(3,7)$, $(5,7)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(2,3,5)$, $(2,3,6)$, $(2,4,7)$, $(3,4,5)$, $(5,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{5}$[21, 6]
$5$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(4,1)$, $(4,5)$, $(4,6)$, $(5,1)$, $(5,6)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,5)$, $(4,5,7)$
$4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
$5$, $6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(3,6)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,3,7)$, $(1,4,7)$, $(2,3,7)$, $(2,4,5)$, $(3,4,6)$, $(5,6,7)$ $\operatorname{Fib}^{2}\boxtimes\operatorname{Sem}$[21]
$1$, $5$, $6$ $(1,5)$, $(1,6)$, $(5,1)$, $(5,6)$, $(6,1)$, $(6,5)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
$1$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,7)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$, $(5,6,7)$
$1$, $4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
$1$, $4$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$4$, $5$, $6$ $(1,4)$, $(1,5)$, $(2,4)$, $(2,5)$, $(4,5)$, $(4,6)$, $(5,4)$, $(5,6)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,5)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,4,6)$, $(1,5,7)$, $(1,6,7)$, $(2,4,7)$, $(2,5,6)$, $(2,6,7)$, $(3,4,5)$, $(3,6,7)$, $(4,5,7)$, $(4,6,7)$, $(5,6,7)$
$1$, $4$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $3$, $5$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(3,1)$, $(3,4)$, $(3,5)$, $(3,7)$, $(4,1)$, $(4,3)$, $(4,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,6)$, $(3,4,7)$, $(3,5,6)$, $(3,6,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
$1$, $4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,6,7)$, $(4,6,7)$, $(5,6,7)$
$1$, $5$, $6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
$4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,1)$, $(5,2)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,5,7)$, $(2,3,7)$, $(2,4,5)$, $(2,5,6)$, $(3,4,5)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
$1$, $2$, $3$, $5$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,6)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,5)$, $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(2,3,5)$, $(2,6,7)$, $(3,4,5)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(5,6,7)$
$1$, $2$, $4$, $6$, $7$ $(1,4)$, $(4,1)$, $(5,2)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,6)$ $(1,2,6)$, $(1,3,4)$, $(1,5,7)$, $(1,6,7)$, $(2,3,5)$, $(2,4,7)$, $(3,6,7)$, $(4,5,6)$, $(4,6,7)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{6}$ [21, 6]
$1$, $2$, $3$, $4$, $7$ $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $4$, $5$, $6$, $7$ $(1,3)$, $(1,5)$, $(2,4)$, $(2,5)$, $(2,6)$, $(3,5)$, $(5,1)$, $(6,1)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,4,7)$, $(1,6,7)$, $(2,3,4)$, $(2,3,6)$, $(2,5,7)$, $(2,6,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{5}$[21, 6]
$1$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,7)$, $(3,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,5)$ $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,6,7)$, $(2,3,5)$, $(2,3,7)$, $(2,4,6)$, $(2,5,6)$, $(3,4,6)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
$1$, $2$, $3$, $4$, $5$, $7$ $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $2$, $3$, $4$, $5$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,6)$, $(5,4)$, $(5,7)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,3,7)$, $(2,3,4)$, $(2,3,5)$, $(2,3,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $3$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
[…]

### v2

@@ 11 $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
77 $(1,7)$, $(2,7)$, $(3,7)$ $(1,2,3)$, $(1,2,4)$, $(1,3,5)$, $(1,6,7)$, $(2,3,6)$, $(2,5,7)$, $(3,4,7)$, $(4,5,6)$ $\operatorname{Fib}\boxtimes\operatorname{Sem}^{2}$[24]
77 $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(6,3)$, $(6,7)$, $(7,3)$ $(1,2,3)$, $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(3,4,5)$, $(4,6,7)$, $(5,6,7)$
44 $(2,4)$, $(3,1)$, $(3,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,4)$, $(2,3,7)$, $(2,5,6)$, $(3,4,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSO}(5)_{6}$[24, 8]
55 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,6)$, $(2,7)$, $(4,3)$, $(4,6)$, $(5,3)$, $(6,3)$, $(6,7)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
66, 77 $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(6,7)$, $(7,6)$ $(1,2,3)$, $(1,2,6)$, $(1,2,7)$, $(1,4,6)$, $(1,5,7)$, $(2,4,7)$, $(2,5,6)$, $(3,4,5)$, $(3,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{6}$[24, 8]
66, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
66, 77 $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(3,7)$, $(5,7)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(2,3,5)$, $(2,3,6)$, $(2,4,7)$, $(3,4,5)$, $(5,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{5}$[24, 8]
55, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(4,1)$, $(4,5)$, $(4,6)$, $(5,1)$, $(5,6)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,5)$, $(4,5,7)$
44, 77 $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
55, 66, 77 $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(3,6)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,3,7)$, $(1,4,7)$, $(2,3,7)$, $(2,4,5)$, $(3,4,6)$, $(5,6,7)$ $\operatorname{Fib}^{2}\boxtimes\operatorname{Sem}$[24]
11, 55, 66 $(1,5)$, $(1,6)$, $(5,1)$, $(5,6)$, $(6,1)$, $(6,5)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
11, 66, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
11, 66, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,7)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$, $(5,6,7)$
11, 44, 77 $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
11, 44, 77 $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
44, 55, 66 $(1,4)$, $(1,5)$, $(2,4)$, $(2,5)$, $(4,5)$, $(4,6)$, $(5,4)$, $(5,6)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,5)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,4,6)$, $(1,5,7)$, $(1,6,7)$, $(2,4,7)$, $(2,5,6)$, $(2,6,7)$, $(3,4,5)$, $(3,6,7)$, $(4,5,7)$, $(4,6,7)$, $(5,6,7)$
11, 44, 77 $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
11, 33, 55 $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(3,1)$, $(3,4)$, $(3,5)$, $(3,7)$, $(4,1)$, $(4,3)$, $(4,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,6)$, $(3,4,7)$, $(3,5,6)$, $(3,6,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
11, 44, 77 $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
11, 22, 33, 44 $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,6,7)$, $(4,6,7)$, $(5,6,7)$
11, 55, 66, 77 $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
44, 55, 66, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,1)$, $(5,2)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,5,7)$, $(2,3,7)$, $(2,4,5)$, $(2,5,6)$, $(3,4,5)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
11, 22, 33, 55 $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,6)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,5)$, $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(2,3,5)$, $(2,6,7)$, $(3,4,5)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(5,6,7)$
11, 22, 44, 66, 77 $(1,4)$, $(4,1)$, $(5,2)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,6)$ $(1,2,6)$, $(1,3,4)$, $(1,5,7)$, $(1,6,7)$, $(2,3,5)$, $(2,4,7)$, $(3,6,7)$, $(4,5,6)$, $(4,6,7)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{6}$ [24, 8]
11, 22, 33, 44, 77 $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
11, 44, 55, 66, 77 $(1,3)$, $(1,5)$, $(2,4)$, $(2,5)$, $(2,6)$, $(3,5)$, $(5,1)$, $(6,1)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,4,7)$, $(1,6,7)$, $(2,3,4)$, $(2,3,6)$, $(2,5,7)$, $(2,6,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{5}$[24, 8]
11, 44, 55, 66, 77 $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,7)$, $(3,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,5)$ $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,6,7)$, $(2,3,5)$, $(2,3,7)$, $(2,4,6)$, $(2,5,6)$, $(3,4,6)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
11, 22, 33, 44, 55, 77 $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
11, 22, 33, 44, 55, 77 $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,6)$, $(5,4)$, $(5,7)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,3,7)$, $(2,3,4)$, $(2,3,5)$, $(2,3,7)$, $(4,5,7)$, $(5,6,7)$
11, 33, 44, 55, 66, 77 $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
[…]